OpenStudy (anonymous):

The quadratic 2x2 + 6x + 3 has what type of roots? Answer Two Irrational Two Rational Two Complex One Rational Double which one is the answer ?

OpenStudy (anonymous):

evaluate the discriminant. b^2-4ac http://en.wikipedia.org/wiki/Discriminant

OpenStudy (anonymous):

okay so its two complex

OpenStudy (anonymous):

b^-4ac = (6)^2 - 4*(2)*(3) = 36-24 = 12 two real roots

OpenStudy (anonymous):

two irrational

OpenStudy (anonymous):

okay thanks